OpenStudy (anonymous):

Solve the following system of equations using elimination: x+y+z=3 x-y+2z=-1 4x+y+z=15

OpenStudy (campbell_st):

well start by reducing the problem to 2 equations with 2 unknowns add the 1st and 2nd equations to eliminate y x + y + z = 3 x -y + 2z = -1 ------------------- 2x +3z = 2 next add the 2nd and 3rd equations to eliminate y x - y + 2z = -1 4x + y + z = 15 ----------------- 5x +3z = 14 so you now have 2 equation in 2 unknowns 2x + 3z = 2 5x + 3z = 14 subtract the 2nd equation from the 1st.. 2x + 3z = 2 - 5x + 3x = 14 ------------------- -3x = -12 now solve for x substitute x into either of the equations above to find z then substitute your values for x and z into any of the original equations to find y. hope this helps

OpenStudy (anonymous):

am i substituting x into one of the original three equations?